A linear model for a set of data was y= -20.6 + 9.72x and produced the following residual plot: [Xscl=1, Yscl=5]
a) Predict the value of y when x= 7.5.
b) Determine the residual if x= 5.

Answer by stanbon(75887)

A linear model for a set of data was y= -20.6 + 9.72x and produced the following residual plot: [Xscl=1, Yscl=5]
a) Predict the value of y when x= 7.5.
f(7.5) = -20.6 + 9.72*7.5 = 52.3
Note: That is the predicted value for y when x = 7.5
--------

b) Determine the residual if x= 5.
Need to know the predicted value f(5) and the actual point (5,?)
=====
Cheers,
Stan H.
==============